# Neurosymbolic Association Rule Mining from Tabular Data

### v1

@@ 1 Introduction
Association Rule Mining (ARM) is a knowledge discovery task that aims to mine commonalities among features of a given dataset as logical implications (Agrawal et al., 1994). It has a plethora of applications in various domains including healthcare (Zhou et al., 2020), recommendation systems (Roy and Dutta, 2022), and anomaly detection (Sarno et al., 2020). Beyond knowledge discovery, ARM plays a crucial role in rule-based interpretable Machine Learning (ML) models such as rule list classifiers (Angelino et al., 2018), particularly in high-stakes decision-making (Rudin, 2019). Such models construct interpretable predictive models using pre-mined rules from ARM algorithms and class labels (Liu et al., 1998; Letham et al., 2015; Angelino et al., 2018). In this paper, we focus on ARM applied to tabular data, a common area of study in ARM research (Kaushik et al., 2023).
@@
To address this research problem, we make the following contributions: i) a novel neurosymbolic ARM method - Aerial+ (Section 3); ii) two comprehensive evaluations of Aerial+ (Section 4) on 5 real-world tabular datasets (Kelly et al., 2023) demonstrating Aerial+’s superiority over seven baselines in knowledge discovery and downstream classification tasks.
Aerial+ is orthogonal to many existing solutions and can be fully integrated with them. It consists of two main steps: the creation of a neural representation of the data using an under-complete denoising autoencoder (Vincent et al., 2008) which captures the associations between features, and the extraction of rules from the neural representation by exploiting the reconstruction mechanism of the autoencoder. The first evaluation uses rule quality, the standard method in ARM literature (Telikani et al., 2020; Kaushik et al., 2023), which shows that Aerial+ can learn a more concise set of high-quality rules than the state-of-the-art with full data coverage. While prior work on ARM predominantly evaluates rule quality, we further evaluate Aerial+ on downstream classification tasks, as part of popular rule-based interpretable ML models such as CORELS (Angelino et al., 2018). The results show that the smaller number of rules learned by Aerial+ leads to faster execution times with similar or higher accuracy. These findings indicate that Aerial+, a neurosymbolic approach, can effectively address the rule explosion problem in ARM research.
@@ 2 Related Work
Rule explosion. ARM suffers from high data dimensionality, leading to excessive rules and long execution times (Telikani et al., 2020; Kaushik et al., 2023). A common remedy is to run ARM with item constraints (Srikant et al., 1997) that focuses on mining rules for the items of interest rather than all (Baralis et al., 2012; Shabtay et al., 2021). Closed itemset mining (Zaki and Hsiao, 2002), another family of solutions, further reduces rule redundancy by identifying only frequent itemsets without frequent supersets of equal support. Another solution is to mine top-k rules based on a given rule quality criteria aiming to control the number of rules to be mined (Fournier-Viger et al., 2012). These methods optimize ARM by reducing search space and improving execution times by limiting rule generation.
Numerical ARM. Another aspect of ARM is its application to numerical data, where many approaches leverage nature-inspired optimization algorithms (Kaushik et al., 2023). Numerical ARM methods aim to find feature intervals that produce high-quality rules based on predefined fitness functions combining rule quality criteria. Notable methods include Bat Algorithm (Yang, 2010; Heraguemi et al., 2015), Grey Wolf Optimizer (Yildirim and Alatas, 2021; Mirjalili et al., 2014), Sine Cosine Algorithm (Mirjalili, 2016; Altay and Alatas, 2021), and Fish School Search (Bastos Filho et al., 2008; Bharathi and Krishnakumari, 2014). Fister et al. (2018) extend numerical ARM methods to categorical data. While numerical ARM methods do not primarily aim to address the rule explosion problem, we include these methods for completeness, as they may yield fewer rules than exhaustive approaches.
Interpratable ML. Besides knowledge discovery, ARM is widely used in rule-based interpretable ML models, which is the standard approach to high-stake decision-making (Rudin, 2019). Examples include associative classifiers such as CBA (Liu et al., 1998), rule set, and rule list learners (Letham et al., 2015; Angelino et al., 2018) that construct rule-based classifiers from pre-mined rules or frequent itemsets via ARM and class labels. Since these models rely on ARM to pre-mine rules, the excessive number of rules and long execution times carry over to downstream interpretable ML tasks, further increasing computational costs. All these methods work with exhaustive ARM approaches such as the popular FP-Growth (Han et al., 2000) to pre-mine frequent itemsets and rules. Numerous versions of FP-Growth have also been proposed to alleviate the aforementioned issues such as Guided FP-Growth (Shabtay et al., 2021) for ARM with item constraints, parallel FP-Growth (Li et al., 2008) and FP-Growth on GPU (Jiang and Meng, 2017) for better execution times.
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset and $f_{i}^{1...c_{i}}=\{0,1\}$ represent categories for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}$ represent the number of categories for feature $f_{i}$ and 0 and 1 indicates the absence or the presence of a feature category in a given transaction (standard ARM setup introduced in Section 2). Input to the autoencoder consists of vectors of dimension $\sum_{i=1}^{k}c_{i}$. Next, a random noise $N\sim[-0.5,0.5]$ is added to each feature category $f_{i}^{c_{j}}$ ($1\leq j\leq c_{i}$), with values clipped to $[0,1]$ using $f_{i}^{c_{j}}=\min(1,\max(0,f_{i}^{c_{j}}+N))$.
The noisy input is propagated through an autoencoder of decreasing dimensions (each layer has half the parameters of the previous layer) with one to three layers per encoder and decoder. The number of layers, training epochs, and batch sizes are chosen depending on the dataset dimensions and number of instances. We found that having more than 3 layers did not improve the performance and we trained for one to two epochs. $tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ is chosen as the activation function in hidden layers. After the encoding and decoding, the softmax function $\sigma$ is applied per feature, $\sigma(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{n}e^{z_{j}}}$, such that the values for categories of a feature sums up to 1 (100 % ):
$\sum_{j=1}^{c_{i}}\sigma(f_{i})_{j}=\sum_{j=1}^{c_{i}}\frac{e^{f_{i,j}}}{\sum_{k=1}^{c_{i}}e^{f_{i,k}}}=\frac{\sum_{j=1}^{c_{i}}e^{f_{i,j}}}{\sum_{k=1}^{c_{i}}e^{f_{i,k}}}=1$
Binary cross entropy (BCE) loss is applied per feature and the results are aggregated:
@@
where $p_{i,j}$ refers to $\sigma(f_{i})_{j}$ and $y_{i,j}$ refers to initial noise-free version of $f_{i}^{j}$. Finally, the learning rate is set to $5e^{-3}$. The Adam (Kingma and Ba, 2014) optimizer is used for gradient optimization with a weight decay of $2e^{-8}$.
@@ 3.3 Rule Extraction Stage
Example. Figure 2 illustrates rule extraction. Assume $weather$ and $beverage$ are features with categories {cold, warm} and {tea, coffee, soda} respectively. The first step is to initialize a test vector of size 5 corresponding to 5 possible categories with equal probabilities per feature, $[0.5,0.5,0.33,0.33,0.33]$. Then we mark $weather(warm)$ by assigning 1 to $warm$ and 0 to $cold$, $[1,0,0.33,0.33,0.33]$, and call the resulting vector a test vector.
Assume that after a forward run, $[0.7,0.3,0.04,0.1,0.86]$ is received as the output probabilities. Since the probability of $p_{weather(warm)}=0.7$ is bigger than the given antecedent similarity threshold ($\tau_{a}=0.5$), and $p_{beverage(soda)}=0.86$ probability is higher than the consequent similarity threshold ($\tau_{c}=0.8$), we conclude with $weather(warm)\rightarrow beverage(soda)$.
@@ 8                   $\mathbf{p}\leftarrow AE(\mathbf{v})$;
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[0,1,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder $AE$ with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
Note that since each feature combination $S\in\mathcal{C}$ is processed independently, Algorithm 1 supports parallel execution. All operations use vector representations, enabling efficient GPU execution. A runtime complexity analysis (Appendix A) shows Aerial+ has linear complexity in the number of transactions $n$ (training), $O(n)$, and depends on the number of antecedents ($a$) and features ($k$), $O(k^{a+1})$ (rule extraction). Since $a$ is typically small (2-4), the complexity is polynomial in $k$. In addition to Algorithm 1, we developed two Aerial+ variants for ARM with item constraints and frequent itemset mining and further describe how different ARM variants integrate into Aerial+, as detailed in Appendix C.
@@ 4 Evaluation
Two sets of experiments evaluate Aerial+ thoroughly: (1) rule quality assessment (Section 4.1), a standard method in ARM research (Kaushik et al., 2023), and (2) testing on downstream classification tasks (Section 4.2) providing input rules to interpretable rule-based classifiers commonly used in high-stake decision-making (Rudin, 2019).
Open-source. The source code of Aerial+, all the baselines and datasets are open-source and can be found at: https://anonymous.4open.science/r/aerial-rule-mining-E7B1.
Hardware. All experiments are run on an AMD EPYC 7H12 64-core CPU with 256 GiB memory. No GPUs were used and no parallel executions were conducted.
@@
Datasets. The experiments use five UCI ML (Kelly et al., 2023) datasets, a standard ARM benchmark, as described in Table 1. Numerical features are discretized into 10 intervals using equal-frequency binning  (Foorthuis, 2020) for algorithms requiring discretization.
@@ 4.1 Experimental Setting 1: Execution Time and Rule Quality Evaluation
The goal of this experimental setting is to compare Aerial+ with seven state-of-the-art ARM algorithms which are given in Table 2 together with their parameters. The comparison is based on the standard evaluation criteria in ARM literature: execution time, number of rules, average support and confidence, and data coverage.
Table 2: Aerial+ and baseline algorithms, along with their types and parameters (R = Aerial+ rules, C = Categories).
@@
Challenges in comparison. Comparing different algorithm types is inherently challenging due to their distinct characteristics. Exhaustive methods identify all rules meeting a given support and confidence threshold, while optimization-based approaches operate within a predefined evaluation limit, improving results up to a point. DL-based ARM methods depend on similarity thresholds for rule quality. Given these differences, we made our best effort to compare algorithms fairly and showed the trade-offs under different conditions. Optimization-based methods were implemented using NiaARM (Stupan and Fister, 2022) and NiaPy (Vrbančič et al., 2018), maintaining original parameter settings, while exhaustive methods used Mlxtend (Raschka, 2018). The antecedent length in exhaustive and DL-based methods is fixed at 2 unless stated otherwise (this can’t be controlled for other methods). Additionally, the minimum support threshold for exhaustive methods is set to half the average support of Aerial+ rules to ensure comparable support values, and ARM-AE’s number of rules per consequent (N) is set to Aerial+’s rule count divided by the number of categories to ensure comparable rule counts.
@@
Execution time and number of rules. Figure 3 shows the change in the number of rules (bars, left y-axis) and execution time (lines, right y-axis) for the exhaustive methods as the number of antecedents increases (top figure, min_support=0.05), or the minimum support threshold decreases (bottom, antecedents=2). The results show that exhaustive methods result in a substantially higher number of rules and execution times when run with a higher number of antecedents or lower minimum support thresholds. The number of rules reaches the order of millions for relatively larger datasets (Chess and Spambase) after 3 and 4 antecedents, while the execution time reaches the level of hours.
Table 3 shows that optimization-based ARM requires long evaluations, hence execution times, to yield higher quality rules. However, improvement in rule quality stagnates after 50,000 evaluations. The results are consistent across datasets (see Appendix D). In contrast, Figure 4 shows that as the number of antecedents increases, Aerial+ generates fewer rules and achieves lower execution times than exhaustive methods (Figure 3). It also significantly outperforms optimization-based methods (Table 3), even with 4 antecedents compared to running the optimization-based method for 10,000 evaluations or more. Importantly, the execution time for Aerial+ includes both training and rule extraction.

### v2

@@ 1 Introduction
Association Rule Mining (ARM) is a knowledge discovery task that aims to mine commonalities among features of a given dataset as logical implications (Agrawal et al., 1994). It has a plethora of applications in various domains including healthcare (Zhou et al., 2020), recommendation systems (Roy and Dutta, 2022), and anomaly detection (Sarno et al., 2020). Beyond knowledge discovery, ARM plays a crucial role in rule-based interpretable Machine Learning (ML) models such as rule list classifiers (Angelino et al., 2018), particularly in high-stakes decision-making (Rudin, 2019). Such models construct interpretable predictive models using pre-mined rules from ARM algorithms and class labels (Liu et al., 1998; Letham et al., 2015; Angelino et al., 2018). In this paper, we focus on ARM applied to tabular data, a common area of study in ARM research (Luna et al., 2019).
@@
To address this research problem, we make the following contributions: i) a novel neurosymbolic ARM method - Aerial+ (Section 3); ii) two comprehensive evaluations of Aerial+ (Section 4) on 5 real-world tabular datasets (Kelly et al., 2023) that demonstrate Aerial+’s superiority over seven baselines both in knowledge discovery and downstream classification tasks.
Aerial+ consists of two main steps: the creation of a neural representation of the data using an under-complete denoising autoencoder (Vincent et al., 2008) which captures the associations between features, and the extraction of rules from the neural representation by exploiting the reconstruction mechanism of the autoencoder. The first evaluation uses rule quality, the standard method in ARM literature (Luna et al., 2019; Kaushik et al., 2023), which shows that Aerial+ can learn a more concise set of high-quality rules than the state-of-the-art with full data coverage. While prior work on ARM predominantly evaluates rule quality, we further evaluate Aerial+ on downstream classification tasks, as part of popular rule-based interpretable ML models such as CORELS (Angelino et al., 2018). The results show that the smaller number of rules learned by Aerial+ leads to faster execution times with similar or higher accuracy. These findings indicate that Aerial+, a neurosymbolic approach, can effectively address the rule explosion problem in ARM research.
@@ 2 Related Work
Rule explosion. ARM has a number of algorithms for finding exhaustive solutions, such as FP-Growth (Han et al., 2000), and HMine (Pei et al., 2007). An extensive survey of such algorithms is provided by Luna et al. (2019). However, ARM suffers from high data dimensionality, leading to excessive rules and long execution times (Telikani et al., 2020; Kaushik et al., 2023). A common remedy is to run ARM with item constraints (Srikant et al., 1997) that focuses on mining rules for the items of interest rather than all (Baralis et al., 2012; Shabtay et al., 2021). Closed itemset mining (Zaki and Hsiao, 2002), another family of solutions, further reduces rule redundancy by identifying only frequent itemsets without frequent supersets of equal support. Another solution is to mine top-k rules based on a given rule quality criteria aiming to control the number of rules to be mined (Fournier-Viger et al., 2012). These methods optimize ARM by reducing search space and improving execution times by limiting rule generation. Aerial+ is orthogonal to these methods and can be fully integrated with them.
Numerical ARM. Another aspect of ARM is its application to numerical data, where many approaches leverage nature-inspired optimization algorithms. Kaushik et al. (2023) presents a comprehensive survey of such algorithms. Numerical ARM methods aim to find feature intervals that produce high-quality rules based on predefined fitness functions, combining rule quality criteria. The best performing methods include Bat Algorithm (Yang, 2010; Heraguemi et al., 2015), Grey Wolf Optimizer (Yildirim and Alatas, 2021; Mirjalili et al., 2014), Sine Cosine Algorithm (Mirjalili, 2016; Altay and Alatas, 2021), and Fish School Search (Bastos Filho et al., 2008; Bharathi and Krishnakumari, 2014). Fister et al. (2018) extends numerical ARM methods to categorical data. While numerical ARM methods do not primarily aim to address the rule explosion problem, we include these methods in our evaluation for completeness, as they may yield fewer rules than exhaustive approaches.
Interpretable ML. Besides knowledge discovery, ARM is widely used in rule-based interpretable ML models, which is the standard approach to high-stake decision-making (Rudin, 2019). Examples include associative classifiers such as CBA (Liu et al., 1998), rule sets, and rule list learners (Letham et al., 2015; Angelino et al., 2018) that construct rule-based classifiers from a set of pre-mined rules or frequent itemsets via ARM and class labels. Since these models rely on ARM to pre-mine rules or frequent itemsets, the excessive number of rules and long execution times carry over to downstream interpretable ML tasks, further increasing computational costs. All these methods work with exhaustive ARM approaches such as FP-Growth to pre-mine frequent itemsets and rules. Numerous versions of FP-Growth have also been proposed to alleviate the aforementioned issues such as Guided FP-Growth (Shabtay et al., 2021) for ARM with item constraints, parallel FP-Growth (Li et al., 2008) and FP-Growth on GPU (Jiang and Meng, 2017) for better execution times.
@@ 3.2 Autoencoder Architecture and Training Stage
Let $F=\{f_{1},f_{2},...,f_{k}\}$ be a set of $k$ features in a tabular dataset (e.g. columns in a table) and $f_{i}^{1...c_{i}}$ represent categories (possible values) for feature $f_{i}$ $(1\leq i\leq k)$ where $c_{i}=|f_{i}|$ indicates the number of categories for feature $f_{i}$. To represent such a tabular dataset as an ARM problem as defined in Section 2, we define the full set of items $I$ as consisting of all possible categories of all features $I=\{f_{1}^{1},...,f_{1}^{c_{1}},...,f_{k}^{1},...,f_{k}^{c_{k}}\}$. Each transaction $T$ in the tabular dataset (e.g., a row in a table) is represented as an itemset $T\subset I$, where each feature contributes exactly one item: $\forall i\in\{1,...,k\},\exists!j\in\{1,...,c_{i}\}\text{ such that }f_{i}^{j}\in T$. Using one-hot encoded representation, $f_{i}^{1...c_{i}}=\{0,1\}$, with 0 and 1 indicating the absence or the presence of a feature category in a given transaction.
The input to the autoencoder consists of vectors of dimension $\sum_{i=1}^{k}c_{i}$. Next, a random noise $N\sim[-0.5,0.5]$ is added to each feature category $f_{i}^{j}$ ($1\leq j\leq c_{i}$), with values clipped to $[0,1]$ using $f_{i}^{j}=\min(1,\max(0,f_{i}^{j}+N))$.
This noisy input is propagated through an autoencoder of decreasing dimensions (each layer has half the parameters of the previous layer) with one to three layers per encoder and decoder. The number of layers, training epochs, and batch sizes are chosen depending on the dataset dimensions and number of instances. We found that having more than 3 layers or training for more than two epochs did not improve the performance. $tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ is chosen as the activation function in hidden layers. After the encoding and decoding, the softmax function $\sigma$ is applied per feature, $\sigma(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{n}e^{z_{j}}}$, such that the values for categories of a feature sums up to 1 (100 % ):
$\sum_{j=1}^{c_{i}}\sigma(f_{i}^{j})=\sum_{j=1}^{c_{i}}\frac{e^{f_{i}^{j}}}{\sum_{k=1}^{c_{i}}e^{f_{i}^{k}}}=\frac{\sum_{j=1}^{c_{i}}e^{f_{i}^{j}}}{\sum_{k=1}^{c_{i}}e^{f_{i}^{k}}}=1$
Binary cross-entropy (BCE) loss is applied per feature, and the results are aggregated:
@@
where $p_{i,j}$ refers to $\sigma(f_{i}^{j})$ and $y_{i,j}$ refers to initial noise-free version of $f_{i}^{j}$. Finally, the learning rate is set to $5e^{-3}$. The Adam (Kingma and Ba, 2014) optimizer is used for gradient optimization with a weight decay of $2e^{-8}$.
@@ 8                   $\mathbf{p}\leftarrow AE(\mathbf{v})$;
Example. Figure 2 shows rule extraction. Let $weather$ and $beverage$ be features with categories cold, warm and tea, coffee, soda, respectively. We start with a test vector of size 5, assigning equal probabilities per feature: $[0.5,0.5,0.33,0.33,0.33]$. Then we mark $weather(warm)$ by assigning 1 to $warm$ and 0 to $cold$, $[1,0,0.33,0.33,0.33]$, and call the resulting vector a test vector. Assume that after a forward run, $[0.7,0.3,0.04,0.1,0.86]$ is received as the output probabilities. Since the probability of $p_{weather(warm)}=0.7$ is bigger than the given antecedent similarity threshold ($\tau_{a}=0.5$), and $p_{beverage(soda)}=0.86$ probability is higher than the consequent similarity threshold ($\tau_{c}=0.8$), we conclude with $weather(warm)\rightarrow beverage(soda)$.
Algorithm. The rule extraction process is given in Algorithm 1. Line 1 stores input feature categories into $\mathcal{F}$. Lines 2-14 iterate over the number of antecedents $i$ and line 3 generates an $i$-feature combination $\mathcal{F}$. Lines 4-13 iterate over the feature combinations $\mathcal{C}$. For each combination $S$, line 5 creates a vector with uniform probabilities per feature category, e.g., $[0.5,0.5,0.33,0.33,0.33]$ vector in the example given above. Line 6 creates a set of test vectors where a combination of feature categories in $S$ are marked per test vector. This corresponds to the $[1,0,0.33,0.33,0.33]$ vector in the example where $weather(warm)$ was marked. Lines 7-13 iterate over the test vectors. Line 8 performs a forward run on the trained autoencoder $AE$ with each test vector $v$. Lines 9-11 compare the output probabilities corresponding to $S$ with a given antecedent similarity threshold $\tau_{a}$, and the algorithm continues with high probability $S$ values. Lines 12-13 compare the output probabilities for the categories in $\mathcal{F}$ that are not in $S$ already ($F\setminus S$), with a given consequent similarity threshold $\tau_{c}$ and stores categories with high probability. Finally, line 14 removes the low support categories from $\mathcal{F}$ so they are ignored in line 3 of the next iteration.
Appendix B describes hyperparameter tuning process for the thresholds $\tau_{a}$ and $\tau_{c}$. A formal justification for Aerial+’s use in ARM is given in Appendix F.
Scalability. A runtime complexity analysis (Appendix A) shows that Aerial+ scales linearly with the number of transactions ($n$) during training and polynomially with the number of features ($k$) during rule extraction. In addition to Algorithm 1, two Aerial+ variants for ARM with item constraints and frequent itemset mining are given in Appendix C, with further explanations on more variants. Since each feature combination $S\in\mathcal{C}$ is processed independently, Algorithm 1 is parallelizable. All computations use vectorized operations, enabling efficient GPU acceleration. Extrapolating the execution times in Section 4, Aerial+ can scale up to tens of thousands of features on a laptop (see Hardware below) in a day.
@@ 4 Evaluation
Two sets of experiments evaluate Aerial+ thoroughly: (1) rule quality assessment (Section 4.1), a standard method in ARM research (Luna et al., 2019), and (2) testing on downstream classification tasks (Section 4.2) providing input rules to interpretable rule-based classifiers commonly used in high-stake decision-making (Rudin, 2019).
@@
Hardware. All experiments are run on an AMD EPYC 7H12 64-core CPU with 256 GiB memory. No GPUs were used, and no parallel execution was conducted.
Datasets. The experiments use five UCI ML (Kelly et al., 2023) datasets (Table 1), a standard ARM benchmark. Numerical features are discretized into 10 intervals via equal-frequency binning (Foorthuis, 2020) for algorithms that require it.
Open-source. The source code of Aerial+, all the baselines and datasets are open-source and can be found at: https://github.com/DiTEC-project/aerial-rule-mining, and Aerial+’s Python package at: https://github.com/DiTEC-project/pyaerial.
@@ 4.1 Experimental Setting 1: Execution Time and Rule Quality Evaluation
The goal of this experimental setting is to compare Aerial+ with seven state-of-the-art ARM algorithms which are given in Table 2 together with their parameters. The comparison is based on the standard evaluation criteria in ARM literature: execution time, number of rules, average support and confidence per rule, and data coverage for the whole dataset.
Challenges in comparison. Comparing different algorithm types is inherently challenging due to their distinct characteristics. Exhaustive methods identify all rules meeting a given support and confidence threshold, while optimization-based approaches operate within a predefined evaluation limit, improving results up to a point. DL-based ARM methods depend on similarity thresholds for rule quality. Given these differences, we made our best effort to compare algorithms fairly and showed the trade-offs under different conditions. Optimization-based methods were implemented using NiaARM (Stupan and Fister, 2022) and NiaPy (Vrbančič et al., 2018), with original parameters, while exhaustive methods used Mlxtend (Raschka, 2018). Antecedent length in exhaustive and DL-based methods is fixed at 2 unless stated otherwise (not controllable in others). The minimum support threshold for exhaustive methods is set to half the average support of Aerial+ rules for comparable support values, and ARM-AE’s number of rules per consequent (N) is set to Aerial+’s rule count divided by the number of categories to ensure comparable rule counts.
Table 2: Aerial+ and baselines with their parameters (R = Aerial+ rules, C = categories).
@@
Execution time and number of rules. Figure 3 shows changes in rule count (bars, left y-axis) and execution time (lines, right y-axis) for exhaustive methods as the number of antecedents increases (top, min_support=0.05) or minimum support decreases (bottom, antecedents=2). Results show that exhaustive methods produce a substantially higher number of rules and has longer execution times with more antecedents or lower support thresholds. Rule counts reach millions for relatively larger datasets (Chess, Spambase) after 3–4 antecedents, while execution time reaches hours.
@@
Table 3 shows that optimization-based ARM requires long evaluations, hence execution times, to yield higher quality rules. However, improvement in rule quality stagnates after 50,000 evaluations. The results are consistent across datasets (see Appendix D). In contrast, Figure 4 shows that as the number of antecedents increases, Aerial+ generates fewer rules and achieves lower execution times than exhaustive methods (Figure 3). It also significantly outperforms optimization-based methods (Table 3), even with 4 antecedents compared to running the optimization-based method for 10,000 evaluations or more. Note that the execution time for Aerial+ includes both training and rule extraction.